OpenStudy (hannahc234):

I will fan and medal whoever can help me with these three questions! Solve this system using substitution 1. 8x+2y=-2 Y=-5x+1 Solve this system using elimination. 2. 4x+y=8 -3x-y=0 3. 3x+2y=-10 2x-5y=3

OpenStudy (maurticewilliams):

what u need help wit

OpenStudy (benlindquist):

easy

OpenStudy (maurticewilliams):

oh let me u help u

OpenStudy (maurticewilliams):

let me get paper

OpenStudy (hannahc234):

Actually I just got the answers lol but thanks anyways.

OpenStudy (maurticewilliams):

hold on im almost finish im gonna giv u the answer just let me work at it no homo

OpenStudy (maurticewilliams):

what was it

OpenStudy (maurticewilliams):

can u tell me

OpenStudy (maurticewilliams):

oh thanks hanna

OpenStudy (hannahc234):

1- x=2 Y=-9 (2,-9) 2-x=8 Y=-24 (8,-24) And I'm still trying to work on 3